Find the slope of the line passing through each pair of points if possible.
P(3,7); Q (6,16)

Answer by mananth(16946)

P(3,7); Q (6,16)
slope m= y2-y1 / x2-x1
m=(16-7)/(6-3)
m=9/3
m=3
...
m.ananth@hotmail.ca